write the equations for 2 lines that intersect at the point (1,0). explain how you know the answer is correct.
Answer by AnlytcPhil(1277)

the equations for 2 lines that intersect at the point (1,0). 

Equations of any line can all be written Ax + By = C
Since you know it goes through the point (1, 0), replace x by 1 and y by 0:
A(1) + B(0) = C
A = C
Now choose any number in the world at random for A and C to equal, and any
number in the world at random for B. I will arbitrarily choose 3 for both
A and C and -2 for B. So Ax + By = C becomes
3x - 2y = 3
Now do it again, but don't make the same choices as before. This time, I
will arbitrarily choose 4 for both A and C and 7 for B. So Ax + By = C
becomes
4x + 7y = 4
explain how you know the answer is correct.
I will check the system
3x - 2y = 3
4x + 7y = 4
to see if (x, y) = (1, 0) is the solution:
3(1) - 2(0) = 3
4(1) + 7(0) = 4
So I get 3 = 3 in the first and 4 = 4 in the second, so it is correct.
Edwin
AnlytcPhil@aol.com